# Laser pulse focusing and energetic electron generation by magnetized plasma lens

### v1

@@ Abstract
An efficient mechanism of laser pulse focusing with the help of shaped underdense plasma target immersed in inhomogeneous magnetic field has been demonstrated. These studies have been carried out with the help of 2-D Particle-In-Cell (PIC) simulation employing the OSIRIS 4.0 platform. It is shown that the divergent magnetic field profile compresses the EM wave pulse in the transverse direction. A comparative investigation with plane and lens shaped plasma geometries has also been conducted to find an optimal configuration for focusing the laser at the desirable location. Furthermore, it is also demonstrated that when the electron cyclotron resonance (ECR) layer is placed at a suitable location where the laser is focused, a highly energetic electron beam gets generated.
@@
EM-wave plasma interaction, Self-focusing, electron beam
@@
Intensity ($I_{0}$) $a_{0}=0.08$ $4.34\times 10^{6}W\ cm^{-2}$ $1.33\times 10^{15}W\ cm^{-2}$
@@
I Introduction
The guiding and focusing of intense electromagnetic waves in plasma channels Ehrlich et al. (1996); Hubbard et al. (2002) have a wide range of applications in fusion plasma, harmonic generation, and laser-plasma accelerators Kaw (2017); Das (2020); Nishida et al. (1987); Ganguli et al. (2016); Gibbon and Förster (1996); Pukhov et al. (1999). Many of these applications require tight focusing of high-intensity lasers. In vacuum, a high-intensity laser can propagate up to a few Rayleigh lengths $(X_{R}=\pi r_{0}^{2}/\lambda)$ without diffraction, where $\lambda$ is the laser wavelength and $r_{0}$ is the spot size of the laser. It is well known that self-focusing in underdense plasma occurs when a relativistically intense laser pulse propagates through it Sprangle et al. (1987). For a sufficiently high-power laser, $P$geq $P_{N}$, the laser becomes self-focused in the plasma medium. Here, $P_{N}$ is the critical power required for nonlinear self-focusing. Physically, ponderomotive force of a focused EM wave pushes the electrons out of high intensity region. The change in density modifies the refractive index of the medium, resulting in the self-focusing of the high-intensity laser. However, this process becomes inefficient if the laser intensity is in the nonrelativistic regime or if the laser power is less than the critical power. Relativistic self-focusing over longer distances can be improved by choosing a preformed density profile. To overcome the difficulty of periodic focusing/defocusing, a localized upward plasma density ramp has been proposed Gupta et al. (2007a, b). Another interesting mechanism to probe the multipetawatt regime is possible through Raman amplification in plasma through the interaction of a short probe pulse with counterpropagating long pump laser beam and longitudinal plasma wave Trines et al. (2011); Malkin et al. (2000). Recent work on long-chirped pulse compression using a preformed density ramp has also suggested the possibility of reaching exawatt or zetawatt levels Hur et al. (2023). At these intensities, the laser pulse will be self-focused. However, in a preformed plasma density profile, there are energy losses resulting from Raman backscatter (RBS) at the quarter-critical density $(n_{c}/4)$ Hur et al. (2005, 2023). These energy loss could be minimized by choosing a sharp density cut-off at $(n_{c}/4)$. Thus, the conventional self-focusing of laser pulse using a plasma medium has limitations, such as the requirement of very high intensity and energy losses due to parametric processes.
The interaction of intense laser pulses with magnetized plasma has recently attracted considerable interest. The laser frequencies being high, the requirement for a magnetic field to elicit a magnetized response from plasma at the laser frequency is quite high. However, for the low frequency CO2 lasers a magnetic field of the order of $\sim kT$ suffices. It is now technologically possible to generate magnetic fields of this strength in laboratory environments Nakamura et al. (2018). Similarly, with the availability of high power microwave pulses (HPMs $P>1GW$) Nevins et al. (1987); Allen et al. (1994); Xiao et al. (2020) interaction of Electromagnetic (EM) waves with magnetized plasma can be explored in the nonlinear regime. Microwave frequencies are in the $\sim$ GHz range, which requires $\sim 0.1-1T$ order magnetic field to elicit a magnetized response from electrons. In simulations, both these regimes can be explored in terms of normalized parameters of the plasma and the EM source. There are three possible distinct geometries termed as the the RL, X, O-mode configurations. In the RL mode the EM wave propagation is parallel to the applied magnetic field, in X and O modes it is perpendicular. While in X mode the laser magnetic field is parallel to applied field in O mode it is perpendicular to the same. Swanson (2020); Stix (1992); Chen et al. (1984). Lately, a number of simulation studies have been carried out for these configurations addressing a wide variety of issues. For instance, study of harmonic generation Dhalia et al. (2023); Maity et al. (2021), localized energy absorption Maity et al. (2022); Vashistha et al. (2022); Juneja et al. (2024); Dhalia et al. (2024), ion heating Juneja et al. (2023); Vashistha et al. (2020), parametric excitation Goswami et al. (2022) and magnetic transparency of plasma medium Mandal et al. (2021) and controlled electron beam generation Choudhary et al. (2023) have been studied.
In this work, we study the role of magnetic field and shaped plasma targets in self-focusing of electromagnetic waves using Particle - In - Cell simulations. Analytical investigation of the role of homogeneous magnetic field on the self-focusing of an intense laser beam in the presence of a transverse as well as axial magnetic field has been carried out by Jha et. al., Jha et al. (2006, 2007). Under a constant applied external magnetic field along the propagation axis, self-focusing of the circularly polarized wave is reported in Molavi Choobini and Ghaffari-Oskooei (2024). A constant magnetic field can induce self-focusing at the discontinuity of the refractive index, which will occur at the vacuum plasma interface of a homogeneous plasma. However, the magnetic field may not necessarily be constant throughout the plasma. We explore the impact of the spatial variation of the magnetic field on self-focusing. For this purpose, a two-dimensional Particle-in-cell simulation was carried out. Through our simulations, we illustrate a novel mechanism that enables the self-focusing of a high-intensity electromagnetic pulse in the presence of an axially diverging magnetic field. We have chosen a cold, homogeneous, fully ionized underdense plasma for our study. In addition, we have also studied the impact of the plasma target shaping. Both a slab and a plasma target in the form of a convex lens has been considered.
The paper is organized as follows. Section II describes the simulation geometry and the choice of parameters. In section III we carry out a comparative study of self-focusing in three different configurations: a) unmagnetized lens, b) magnetized lens c) magnetized slab. Subsection III.1 presents analytical calculations for spot size evolution in the chosen geometry and describes the effect of polarization on self-focusing. In section IV the energetic electron beam generation has been demonstrated. In section V we summarize our findings and conclude.
@@ I Introduction
II Simulation Details
[…]
The chosen external magnetic field profile has a divergent nature from the center line $y=110l_{N}$, which satisfies the required condition of .$\vec{\nabla}.\vec{B}=0$. Here, the value of parameter $(\delta=0.005)$ is chosen such that the associated gyrofrequency $(\omega_{ce})$ due to the external magnetic field at the end of the target plasma stays larger than the incident EM wave frequency $[\omega_{ce}(x=160l_{N})>\omega_{EM}]$. This ensures that the group velocity of the EM wave, within the plasma stays finite. The simulation run time was $300\omega_{pe}^{-1}$. During this time, ion response is negligible compared to electrons and they merely act as a neutralizing background Dhalia et al. (2024); Maity et al. (2022). Thus, the driving source of focusing for incoming EM wave pulse is associated with the dynamics of electrons, which being lighter, respond at such a fast time scale. However, it should be noted that both the electron and ion dynamics have been considered in simulation. The normalized vector potential of EM wave is taken $a_{0}=0.08$ to stay within a non-relativistic regime. This intensity has been chosen such that the self-focusing due to relativistic nonlinearity can be avoided and we can have a direct comparison of unmagnetized and magnetized plasma.
The chosen normalized parameters can represent both a laser-plasma interaction as well as also microwave-plasma interaction by using appropriate normalization factors. This has been illustrated in Table -1.
The microwave system features a pulse, high power microwave of spot size 68.56 cm and frequency 6.7 GHz and power of approximately 4 GW incident on a plasma lens of height 171.4 cm and width of 85.7 cm Nevins et al. (1987); Allen et al. (1994). The external magnetic field required here is of order $\sim 0.44T$ or 4.4 kilo gauss, which can be generated through a simple permanent magnet or an electromagnet. Similarly, a laser system will feature a high power PW, $\text{CO}_{2}$ laser of spotsize 63.2 $\mu m$ , incident on a plasma lens of height $158\mu m$ and width of $79\mu m$. The external magnetic field required is of order $4.71kT$ or 47.1 mega gauss Choudhary et al. (2025); Hao et al. (2025).
Figure 9: Figure demonstrates the plasma lens situated in a high gradient external magnetic field. The ECR point represent the location at plasma edge where EM wave frequency matches with electron cyclotron frequency.
III Magnetic field induced focusing
@@
These studies thus illustrate that both the shape of plasma and the profile of the magnetic field play crucial roles in EM wave pulse focusing. It should be noted that this is a novel approach. The focusing here does not require relativistic intensities of EM wave-pulse or preformed density profiles. In fact even a moderately intense pulse can be focused at a desired location. Another important feature of this technique is that almost no energy is dissipated into plasma during this interaction. Figure 6 shows time evolution of total energy, field energy and kinetic energy for the case of magnetic lens. From this plot we observe that the total energy remains almost constant. There is a conversion from field to kinetic energy in the beginning of the interaction. Thereafter, however, the kinetic energy again converts back into field energy. We now try to understand the contrasting response of focusing displayed by the LCP and the RCP EM waves in the next subsection.
III.1 Interpretation
We have also carried out simulations for an incident wave with a definite RCP and LCP form of polarization. Here the waves do not break up in two pulses as expected. However, only the RCP wave shows focusing whereas the wave pulse with LCP shows divergence as seen from figure 7.
This distinctive behavior of focal spot evolution for the two polarizations have been understood analytically by applying a local approximation to the theory put forth for the homogeneous magnetic field by Jha et al Jha et al. (2006, 2007). In there study the source dependent expansion (SDE) method Esarey et al. (1997), is used to study the evolution of the laser spot. The expression for laser spot evolution in the presence of an axial magnetic field is given as Jha et al. (2007),
$\frac{r_{s}^{2}}{r_{0}^{2}}=1+(1-\frac{P}{P_{c}})\frac{x^{2}}{X_{R}^{2}}$ (2)
@@ III.1 Interpretation
$S=\frac{\omega_{EM}^{4}(\omega_{EM}+\sigma\omega_{ce})^{4}}{(\omega_{EM}^{2}-\omega_{ce}^{2})^{4}}$ (3)
Here, $\omega_{ce}=eB_{0}/m$. $X_{R}=(k_{0}^{2}r_{0}^{2}/2)$ is the Rayleigh length. We have plotted the critical power $P_{c}$ with the help of equation (3) along the axis for the changing magnetic field in our system. The variation of the critical power with distance $x$ has been shown in figure 8$(a)$. It should be noted from the figure that the value of critical power is very high for the LCP wave, whereas for the RCP wave it is quite low. The power of our chosen EM wave pulse exceeds the critical power for the RCP wave and hence focusing is observed.
@@
IV Magnetic field induced focused electron beam generation
In the previous section, it has been shown that the RCP wave gets efficiently focussed with the help of a magnetized plasma lens. We show here that by appropriately tailoring the magnetic field profile with the help of the parameter $\delta$ we can produce energetic electron beam. This happens when the magnetic field profile is tailored to have the ECR resonance lie within the plasma. In our studies we achieve this by choosing the following form of the magnetic field profile with the choice of $\delta=0.02$.

### v2

@@ Abstract
An efficient novel mechanism of laser pulse focusing with the help of a shaped underdense plasma target immersed in an inhomogeneous magnetic field has been demonstrated. These studies have been carried out with the help of 2-D Particle-In-Cell (PIC) simulation employing the OSIRIS 4.0 platform. It is shown that the divergent magnetic field profile compresses the EM wave pulse in the transverse direction. A comparative investigation with plane and lens-shaped plasma geometries has also been conducted to find an optimal configuration for focusing the laser at the desirable location. Furthermore, it is also demonstrated that when the electron cyclotron resonance (ECR) layer is placed at a suitable location where the laser is focused, a highly energetic electron beam gets generated.
@@
Keywords: EM-wave plasma interaction, Magnetic field induced focusing, Electron beam
1 Introduction
The guiding and focusing of intense electromagnetic waves in plasma channels (Ehrlich et al., 1996; Hubbard et al., 2002; Esarey et al., 1997) have a wide range of applications in fusion plasma, harmonic generation, and laser-plasma accelerators (Kaw, 2017; Das, 2020; Nishida et al., 1987; Ganguli et al., 2016; Gibbon & Förster, 1996; Pukhov et al., 1999). Many of these applications require tight focusing of high-intensity lasers. In vacuum, a high-intensity laser can propagate up to a few Rayleigh lengths $(X_{R}=\pi r_{0}^{2}/\lambda)$ without diffraction, where $\lambda$ is the laser wavelength and $r_{0}$ is radius of the spot size of the laser. It is well known that self-focusing in underdense plasma occurs when a relativistically intense laser pulse propagates through it (Sprangle et al., 1987). For a sufficiently high-power laser, $P$geq $P_{N}$, the laser becomes self-focused in the plasma medium. Here $P_{N}$ is the critical power required for nonlinear self-focusing. Physically, the ponderomotive force of a focused EM wave pushes the electrons out of the high-intensity region. The change in density modifies the refractive index of the medium, resulting in the self-focusing of the high-intensity laser. However, this process becomes inefficient if the laser intensity is in the nonrelativistic regime or if the laser power is less than the critical power. Relativistic self-focusing over longer distances can be improved by choosing a preformed density profile. To overcome the difficulty of periodic focusing/defocusing, a localized upward plasma density ramp has been proposed (Gupta et al., 2007b, a). Recent work on long-chirped pulse compression using a preformed density ramp has also suggested the possibility of reaching exawatt or zetawatt levels (Hur et al., 2023). At these intensities, the laser pulse will be self-focused. However, in a preformed plasma density profile, there are energy losses resulting from Raman backscatter (RBS) at the quarter-critical density $(n_{c}/4)$ (Hur et al., 2005, 2023). These energy losses could be minimized by choosing a sharp density cut-off at $(n_{c}/4)$. Thus, the conventional self-focusing of a laser pulse using a plasma medium has limitations, such as the requirement of very high intensity and energy losses due to parametric processes.
The interaction of intense laser pulses with magnetized plasma has recently attracted considerable interest. The laser frequencies being high, the requirement for a magnetic field to elicit a magnetized response from plasma at the laser frequency is quite high. However, for the low-frequency $CO_{2}$ lasers, a magnetic field of the order of $\sim kT$ suffices. It is now technologically possible to generate magnetic fields of this strength in laboratory environments (Nakamura et al., 2018). Similarly, with the availability of high-power microwave pulses (HPMs $P>1GW$) (Nevins et al., 1987; Allen et al., 1994; Xiao et al., 2020) interaction of electromagnetic (EM) waves with magnetized plasma can be explored in the nonlinear regime. Microwave frequencies are in the $\sim$ GHz range, which requires $\sim 0.1-1T$ order magnetic field to elicit a magnetized response from electrons. In simulations, both these regimes can be explored in terms of normalized parameters of the plasma and the EM source. There are three possible distinct geometries termed as the RL, X, and O-mode configurations. In the RL mode the EM wave propagation is parallel to the applied magnetic field; in X and O modes it is perpendicular. While in X mode the laser magnetic field is parallel to the applied field; in O mode it is perpendicular to the same. (Swanson, 2020; Stix, 1992; Chen et al., 1984). Lately, a number of simulation studies have been carried out for these configurations addressing a wide variety of issues. For instance, study of harmonic generation (Dhalia et al., 2023; Maity et al., 2021), localized energy absorption (Maity et al., 2022; Vashistha et al., 2022; Juneja et al., 2024; Dhalia et al., 2024), ion heating (Juneja et al., 2023; Vashistha et al., 2020), parametric excitation (Goswami et al., 2022) and magnetic transparency of plasma medium (Mandal et al., 2021) and controlled electron beam generation (Choudhary et al., 2023) have been studied.
In this work, we study the role of magnetic field and shaped plasma targets in self-focusing of electromagnetic waves using Particle - In - Cell simulations. Analytical investigation of the role of a homogeneous magnetic field on the self-focusing of an intense laser beam in the presence of a transverse as well as axial magnetic field has been carried out by (Jha et al., 2006, 2007). Under a constant applied external magnetic field along the propagation axis, self-focusing of the circularly polarized wave is reported in (Molavi Choobini & Ghaffari-Oskooei, 2024). A constant magnetic field can induce self-focusing at the discontinuity of the refractive index, which will occur at the vacuum plasma interface of a homogeneous plasma. However, in real scenario the magnetic field may not necessarily be constant throughout the plasma. We explore the impact of the spatial variation of the magnetic field on self-focusing. For this purpose, a two-dimensional Particle-in-cell simulation was carried out. Through our simulations, we illustrate a novel mechanism that enables the self-focusing of a high-intensity electromagnetic pulse in the presence of an axially diverging magnetic field. We have chosen a homogeneous, fully ionized underdense plasma for our study. In addition, we have also studied the impact of the plasma target shaping. Both a slab and a plasma target in the form of a convex lens has been considered.
The paper is organized as follows. Section 2 describes the simulation geometry and the choice of parameters. In section 3 we carry out a comparative study of self-focusing in three different configurations: a) magnetized lens, b) unmagnetized lens c) magnetized slab. Subsection 3.1 presents analytical calculations for spot size evolution in the chosen geometry and describes the effect of polarization on self-focusing. In section 4 the energetic electron beam generation has been demonstrated. In section 5 we summarize our findings and conclude.
@@ 1 Introduction
Intensity ($I_{0}$) $a_{0}=0.08$ $4.34\times 10^{6}W\ cm^{-2}$ $9.86\times 10^{13}W\ cm^{-2}$
@@
2 Simulation Details
[…]
$c(y-110)^{2}+60\leq x\leq-c(y-110)^{2}+160$ (1)
Here, parameter $c$ determines shape of the plasma. For the plasma lens geometry $c=0.005$ and $c=0$ will give a plasma slab geometry. The plasma region extends from $x=60l_{N}$ to $x=160l_{N}$ and $y=10l_{N}$ to $y=210l_{N}$ and the central axis is along $y=110l_{N}$. The incident EM wave has frequency $\omega_{EM}=1.2\omega_{pe}$. Here $\omega_{pe}$ represents normalized plasma frequency. The spatial profile of the pulse is Gaussian. The Incident laser pulse focused initially at the vacuum plasma interface at $x=60l_{N}$. The pulse duration of the incoming wave is $50\omega_{pe}^{-1}$ and the transverse initial spot-size has a diameter of $80l_{N}$. The vacuum diffraction length of the incoming pulse is $k_{0}r_{0}^{2}/2=960l_{N}$. Thus within entire simulation box length $(300l_{N})$ incoming pulse’s spot-size remains unchanged until it interacts with plasma. The spot-size initially is taken considerably smaller than the transverse extent $(220l_{N})$ of the plasma to minimize the effect of spherical aberration (Li et al., 2022). The profile of the external magnetic field has the following form,
The chosen external magnetic field profile has a divergent nature from the center line $y=110l_{N}$, which satisfies the required condition of .$\vec{\nabla}.\vec{B}=0$. Here, the value of parameter $(\delta=0.005)$ is chosen such that the associated gyrofrequency $(\omega_{ce})$ due to the external magnetic field at the end of the target plasma stays larger than the incident EM wave frequency $[\omega_{ce}(x=160l_{N})>\omega_{EM}]$ as shown in figure 1(c). This ensures that the group velocity of the EM wave, within the plasma stays finite. The simulation run time was $300\omega_{pe}^{-1}$. During this time, ion response is negligible compared to electrons and they merely act as a neutralizing background (Dhalia et al., 2024; Maity et al., 2022). Thus, the driving source of focusing for incoming EM wave pulse is associated with the dynamics of electrons, which being lighter, respond at such a fast time scale. However, it should be noted that both the electron and ion dynamics have been considered in simulation. The normalized vector potential of EM wave is taken $a_{0}=0.08$ to stay within a non-relativistic regime. This intensity has been chosen such that the self-focusing due to relativistic nonlinearity can be avoided and we can have a direct comparison of unmagnetized and magnetized plasma. The chosen normalized parameters can represent both a laser-plasma interaction as well as also microwave-plasma interaction by using appropriate normalization factors. This has been illustrated in Table -1.
The microwave system features a pulse, high power microwave of spot size 68.56 cm and frequency 6.7 GHz and power of approximately 16 GW incident on a plasma lens of height 171.4 cm and width of 85.7 cm (Nevins et al., 1987; Allen et al., 1994). The external magnetic field required here is of order $\sim 0.44T$ or 4.4 kilo gauss, which can be generated through a simple permanent magnet or an electromagnet. Similarly, a laser system will feature a moderate power 0.03 TW, $\text{CO}_{2}$ laser of spotsize 63.2 $\mu m$ , incident on a plasma lens of height $158\mu m$ and width of $79\mu m$. The external magnetic field required is of order $4.71kT$ or 47.1 mega gauss (Choudhary et al., 2025; Hao et al., 2025).
@@ 2 Simulation Details
3 Magnetic field induced focusing
@@
These studies thus illustrate that both the shape of plasma and the profile of the magnetic field play crucial roles in EM wave pulse focusing. It should be noted that this is a novel approach. The focusing here does not require relativistic intensities of EM wave-pulse. In fact even a moderately intense pulse can be focused at a desired location. Another important feature is that almost no energy is dissipated into plasma during this interaction. Figure 6 shows time evolution of total energy, field energy and kinetic energy for the case of magnetic lens. From this plot we observe that the total energy remains almost constant. There is a conversion from field to kinetic energy in the beginning of the interaction. Thereafter, however, the kinetic energy again converts back into field energy. We now try to understand the contrasting response of focusing displayed by the LCP and the RCP EM waves in the next subsection.
3.1 Interpretation
We have also carried out simulations for an incident wave with a definite RCP and LCP form of polarization. Here the waves do not break up in two pulses as expected. However, only the RCP wave shows focusing whereas the wave pulse with LCP shows divergence as seen from figure 7. This distinctive behavior of focal spot evolution for the two polarizations have been understood analytically by applying a local approximation to the theory put forth for the homogeneous magnetic field by (Jha et al., 2006, 2007). In there study the source dependent expansion (SDE) method (Esarey et al., 1997), is used to study the evolution of the laser spot. The expression for laser spot evolution in the presence of an axial magnetic field is given as ,
$\centering\frac{r_{s}^{2}}{r_{0}^{2}}=1+(1-\frac{P}{P_{c}})\frac{x^{2}}{X_{R}^{2}}\@add@centering$ (3)
@@ 3.1 Interpretation
$\centering S=\frac{\omega_{EM}^{4}(\omega_{EM}+\sigma\omega_{ce})^{4}}{(\omega_{EM}^{2}-\omega_{ce}^{2})^{4}}\@add@centering$ (5)
Here, $\omega_{ce}=eB_{0}/m$. $X_{R}=(k_{0}^{2}r_{0}^{2}/2)$ is the Rayleigh length. We have plotted the critical power $P_{c}$ with the help of equation (4) and (5) along the axis for the changing magnetic field in our system. The variation of the critical power with distance $x$ has been shown in figure 8$(a)$. It should be noted from the figure that the value of critical power is very high for the LCP wave, whereas for the RCP wave it is quite low. The power of our chosen EM wave pulse exceeds the critical power for the RCP wave and hence focusing is observed.
@@
Figure 9: Figure (a) shows the FWHM of incident EM wave for four different curvature of convex plasma geometry and they are fitted with a second order polynomial $a_{2}x^{2}+a_{1}x+a_{0}$. In figure (b) we plot the coefficient $a_{2}$ with respect to curvature parameter $c$.
3.2 Effect of lens curvature on focusing
We established in the preceding section that incident RCP wave pulses are strongly focused under the magnetized plasma lens. Here, we want to demonstrate the effect of plasma lens curvature on magnetic field-induced focusing. For this, we have incident RCP waves on five distinct shapes corresponding to the curvature parameter $(c=0,0.001,0.003,0.005,0.007)$ of plasma and kept the magnetic field profile $(\delta=0.005)$ constant in all five cases. The figure 9(a) shows the evolution of the beam waist $r_{s}$ of an RCP wave pulse through several plasma lenses. This effectively demonstrates that increasing the curvature of the plasma lens improves focusing. The curves fit well on a second-degree polynomial ($a_{2}x^{2}+a_{1}x+a_{0}$) within the plasma boundary. In the case of $(c=0.007)$, a minimal spot is observed within the plasma lens, and spot size decreases in a quadratic pattern. It is also worth noting that for a slab geometry $(c=0)$, focusing is lowest and spot size decreases in a linear fashion. Figure 9(b) demonstrates that as the curvature $c$ grows, so does the coefficient $a_{2}$ in a linear fashion. Obviously, the curvature parameter should not exceed a particular limit so that the lens height is comparable to or less than the laser spot size; otherwise, spherical aberration will impede the focusing. This concludes that the curvature of the plasma lens provides additional focusing on top of magnetic field-induced focusing.
Figure 10: Figure demonstrates the plasma lens $(c=0.005)$ situated in a high gradient external magnetic field. The ECR point represent the location at plasma edge where EM wave frequency matches with electron cyclotron frequency.
4 Magnetic field induced focused electron beam generation
In the previous section, it has been shown that the RCP wave gets efficiently focused with the help of a magnetized plasma lens. We show here that by appropriately tailoring the magnetic field profile with the help of the parameter $\delta$ we can produce energetic electron beam. This happens when the magnetic field profile is tailored to have the ECR resonance lie within the plasma. In our studies we achieve this by choosing the following form of the magnetic field profile with the choice of $\delta=0.02$.